OpenStudy (anonymous):

Which equation will have the same solution as 3 – 5(2x + 1) = 3(3 + x)? A. 11x = –13 B. 11x = 13 C. 13x = –11 D. 13x = 11

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

idk cuz i i dont get any of the anwers

rishavraj (rishavraj):

@rocky1718 did u find the value of x ..by solving 3 – 5(2x + 1) = 3(3 + x)????

OpenStudy (anonymous):

no

rishavraj (rishavraj):

just solve for "x" first...

OpenStudy (anonymous):

3 – 5(2x + 1) = 3(3 + x) 3(10x+5)=(9+3x) 30x+5=9+3x 27x+5+9

OpenStudy (anonymous):

thats how i did it

rishavraj (rishavraj):

@rocky1718 its supposed to be 3 - 10x - 5 = 9 + 3x go further...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so subtract 10 and 3 is 7

rishavraj (rishavraj):

wht u mean... its supposed to be -10x - 2 = 9 + 3x -13x = 11

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx

rishavraj (rishavraj):

@rocky1718 u r most welcome...:)))